OpenStudy (anonymous):

Solve for n: log n + log n log 2 <= n

OpenStudy (anonymous):

Is this base 10 or base e? And, is this the case? \[ \log(n)+\log(n)\cdot\log(2)=n \]

OpenStudy (anonymous):

we can assume it's base e

OpenStudy (anonymous):

Oh, wait, that was a stupid question, never mind, here: \[ \log(n)+\log(n)\cdot\log(2)\leq n \] There's no elementary way to solve this, just a note (not that I can see, at least). But, we know that: \[ \frac{d}{dn}\log(n)(1+\log(2))\le \frac{d}{dn}n, n\geq1+\log(2) \]And (with some quick algebra): \[ \log(n)(1+\log(2))\lt n, n\in (1, 1+\log(2)) \]And it is obvious to see that: \[ \log(n)(1+\log(2))\lt n, n\in (0, 1] \]Thus, we are done, and we see that: \[ n\in(0, 1]\cup(1, 1+\log(2))\cup[1+\log(2), \infty) = (0, \infty) \]Or: \[ n\gt 0 \]